OpenStudy (anonymous):

I am trying to simplify this and I can't get it √(-2,-0)^2+(3,0)^2

hartnn (hartnn):

are you trying to find distance between (-2,0) and (3,0) ??

hartnn (hartnn):

or (-2,3) and (0,0) ??

hartnn (hartnn):

if its (-2,3) and (0,0), then there's a shortcut, but we'll use distance formula only, \(d= \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) so lets first find x1-x2 = -2-0 =... ? y1-y2 = 3-0 = ... ?

OpenStudy (anonymous):

Yes I need the distance between them

OpenStudy (anonymous):

Then simplify under the radical

hartnn (hartnn):

but the points are these, (-2,3) and (0,0), right ? x1-x2 = -2-0 =-2 y1-y2 = 3-0 = 3 so, d= \(\sqrt{(-2)^2+3^2}=..?\)